# Together We Go Further: LLMs and IDE Static Analysis for Extract Method Refactoring

### v1

@@ Abstract.
Excessively long methods that encapsulate multiple responsibilities within a single method are challenging to comprehend, debug, reuse, and maintain. The solution to this problem, a hallmark refactoring called Extract Method, consists of two phases: (i) choosing the statements to extract and (ii) applying the mechanics to perform this refactoring. While the application part has been a staple feature of all modern IDEs, they leave it up to developers to choose the statements to extract. Choosing which statements are profitable to extract has been the subject of many research tools that employ hard-coded rules to optimize software quality metrics. Despite steady improvements, these tools often fail to generate refactorings that align with developers’ preferences and acceptance criteria. Given that Large Language Models (LLMs) have been trained on large code corpora, they are more likely to imitate human behavior. If we harness the familiarity of LLMs with the way developers form functions, we could suggest refactorings that developers are likely to accept.
@@
In this paper, we advance the science and practice of refactoring by augmenting the refactoring capabilities of IDEs with the power of LLMs to perform Extract Method. Our formative study on 2,849 Extract Method scenarios revealed that LLMs are very effective in giving expert suggestions, yet they are unreliable: up to 62.3% of the suggestions are hallucinations. We propose a novel approach that synergistically combines the creative potential of LLMs with static analysis to enhance the refactoring suggestions generated by LLMs. Additionally, we utilize the safety measures of static analysis within IDEs to execute refactorings safely. Starting from candidates suggested by LLMs, we filter, further enhance, and then rank suggestions based on static analysis techniques from program slicing. We designed, implemented, and evaluated this approach in an IntelliJ IDEA plugin called EM-Assist. We empirically evaluated EM-Assist on a diverse, publicly available corpus that other researchers used in the past. The results show that EM-Assist outperforms previous state-of-the-art tools: at 1% tolerance, EM-Assist suggests the correct refactoring among its top-5 suggestions 60.6% of times, compared to 54.2% reported by existing ML models, and 52.2% reported by existing static analysis tools. When we replicated 2,849 actual Extract Method instances from open-source projects, EM-Assist’s recall rate was 42.1% compared to 6.5% for its peers. Furthermore, we conducted firehouse surveys with 20 industrial developers and suggested refactorings on their recent commits. 81.3% of the respondents agreed with the recommendations provided by EM-Assist. This shows the usefulness of our approach and ushers us into a new era of refactoring when LLMs become effective AI assistants for developers.
copyright: none
1. Introduction
@@
Excessively long methods that encapsulate multiple responsibilities within a single method are challenging to comprehend, debug, reuse, and maintain (Tsantalis and Chatzigeorgiou, 2011; Banker et al., 1993; Fowler, 1999). To mitigate these issues, software developers use Extract Method refactoring – a hallmark refactoring (Fowler, 1999) supported by all modern IDEs. This refactoring involves moving a block of statements from a host method to a brand new method, passing the used variables as parameters to the new method, and adding a call to the new method from the host method. This refactoring is reported as being among the top-5 most frequently performed in practice (Negara et al., 2013; Murphy-Hill et al., 2012; Tsantalis et al., 2022) both for manual and automated refactoring.
@@
The process of performing an Extract Method consists of two phases: (i) choosing the statements to extract and (ii) applying the mechanics to perform this refactoring. While the application part has been a staple feature of all modern IDEs, they leave it up to developers to choose the statements to extract. The research community developed various techniques for suggesting these. Some research tools employ static analysis (Maruyama, 2001; Tsantalis and Chatzigeorgiou, 2011; Charalampidou et al., 2017; Yang et al., 2009a; Tiwari and Joshi, 2022), while others employ machine learning (ML) models (Xu et al., 2017; Cui et al., 2023). These tools employ various software quality metrics (e.g., cohesion of statements), which they maximize either based on heuristics or by training ML classifiers.
@@
While these tools excel in adhering to software engineering principles like the Single Responsibility Principle (Martin, 2017), they exhibit relatively low recall rates when evaluated on real-world Extract Method instances. We posit that the decision to refactor is both a science and an art. When refactoring, developers use both their knowledge of software engineering principles and rely on their own experience and subjective interpretation of the code context and what makes a good method. This can be a reason for the developers’ reluctance to use automated refactoring support (Silva et al., 2016) and can explain the large gap between high metrics scores of refactoring recommendation approaches and their low acceptance by developers (Fakhoury et al., 2019; Scalabrino et al., 2019; Dilhara et al., 2021, 2022).
@@
Recently, Large Language Models (LLMs) (Feng et al., 2020; OpenAI, 2023; GoogleAI, 2023) are emerging as powerful companions for software engineering tasks: code completion (Ciniselli et al., 2022), code summarization (Feng et al., 2020), commit message generation (Aleksandra et al., 2023), code transformation (Ramos et al., 2023; Dilhara et al., 2024), reproducing bug reports (Sidong and Chunyang, 2024), etc. Given LLMs’ training on extensive code repositories that contain billions of methods written by actual developers, we hypothesize that they are more likely to imitate human behavior by mimicking how developers form functions, thus are likely to suggest refactorings that developers would accept. Our formative study on 2,849 Extract Method scenarios revealed that LLMs are very effective in giving expert suggestions. They are also very prolific, producing 12,387 suggestions for 2,849 Extract Method scenarios. However, we also discovered that LLMs are unreliable. 62.3% of their suggestions are hallucinations, i.e., they seem plausible at first, but are actually deeply flawed: 45.7% of the suggested refactorings contain code fragments that are invalid to extract (e.g., would produce compile errors), and 16.6% are not useful (e.g., the suggestion includes the whole body of the host method).
@@
To advance the state of the art and practice for refactoring, we aim to bridge several important gaps. First, we bridge the gap between the suggestion of refactorings and developers’ actual practices by grouping statements into methods that resemble human-written code. Second, we bridge the gap between suggesting and applying refactorings by supporting the whole end-to-end process in a way that provides maximum automation while taking into account human input. Thus, we have designed, implemented, and evaluated EM-Assist, an IntelliJ IDEA plugin, that supports Java and Kotlin Extract Method refactorings. EM-Assist synergistically combines (i) the creative capabilities of LLMs, (ii) static analysis techniques to filter and enhance LLM-provided suggestions, and (iii) the full power of a state-of-the-practice commercial IDE to apply refactorings safely. EM-Assist first repeatedly prompts the LLM in a few-shot learning style to generate a diverse range of refactoring suggestions. Subsequently, it eliminates two kinds of hallucinations: (i) it employs static analysis techniques from the IDE to eliminate invalid suggestions (i.e., illegal groupings of code statements), and (ii) it eliminates suggestions that are not useful (e.g., that include the whole body of the host method, or one single line). Then it further enhances the remaining valid suggestions based on program slicing techniques. Following this step, it employs three ranking mechanisms to prioritize high-quality suggestions so that it would not overwhelm the developer with too many suggestions. It then presents the ranked suggestions to the developers. Lastly, it encapsulates the user-chosen candidate into a refactoring command and uses the IDE to correctly execute the refactoring.
@@
We designed a comprehensive empirical evaluation to determine the benefits of our novel approach. First, we determine the effectiveness of several LLMs in generating refactoring recommendations. Our results show that LLMs are prolific, generating on average 20 suggestions per method, yet on average, 11 are hallucinations. Second, we conduct a sensitivity analysis to tune hyper-parameters of the LLMs (e.g., the “temperature” which controls the degree of generations’ variability) and to measure the improvements brought by our enhancements. Third, to determine the effectiveness of EM-Assist, we use a diverse, publicly available corpus (Xu et al., 2017) that other researchers used in the past. Furthermore, we replicated 2,849 of actual Extract Method instances from open-source projects. The results show that EM-Assist outperforms state-of-the-art static-analysis tools such as JDeodorant (Tsantalis and Chatzigeorgiou, 2011), JExtract (Silva et al., 2015), SEMI (Charalampidou et al., 2017), LiveRef  (Fernandes et al., 2022, 2023), and also outperforms machine learning-based techniques like REMS (Cui et al., 2023) and GEMS (Xu et al., 2017). EM-Assist is able to suggest the ground truth extract method refactoring among the top-5 candidates 60.6% of times, compared to 54.2% reported by existing ML models, and 52.2% reported by existing static analysis tools.
@@
Moreover, to assess the practicality and effectiveness of refactoring recommendations generated by EM-Assist, we employed firehouse surveys (Murphy-Hill et al., 2015) with 16 industrial developers from a reputable software company. We contacted 20 developers and presented them with refactoring suggestions for lengthy methods they previously committed into their software repository. 80% of the developers responded, with 81.3% of respondents agreeing that the suggestions were beneficial and suitable for application in their codebase. They wish to see EM-Assist in production in IntelliJ IDEA and use it in their daily coding. This demonstrates that EM-Assist generates suggestions that are highly likely to be embraced by developers and ushers us into a new era of refactoring when developers are accepting AI assistants.
In summary, this paper makes the following key contributions:
@@
    To the best of our knowledge, we present the first approach to synergistically combine the creativity of LLMs with the safety of static analysis and IDEs. Starting from candidates suggested by LLMs, we filter, further enhance, and then rank suggestions based on static analysis. Thus, we effectively bridge the gap between refactoring recommendation and application, and the gap between refactoring suggestions and developer practices.
@@
    We discovered best practices for prompting LLMs to produce effective refactoring suggestions and we reveal LLMs’ strengths and weaknesses for refactoring tasks.
@@
    Our comprehensive empirical evaluation on a corpus used by others, as well as significantly extended corpus that replicates 2,849 real-world refactorings shows that EM-Assist outperforms traditional static analysis-based techniques and machine learning approaches in automatically suggesting and performing Extract Method refactoring. Moreover, our survey with 16 industrial developers provides insights about the reasons why Java developers accept or reject suggestions generated by EM-Assist.
@@
To aid reproducibility in SE research, we make EM-Assist, a demo, and all our experimental data available anonymously (Anonymous, 2023).
2. Motivating Example
@@
Figure 1. The numbered code snippets represent 1: an extract function refactoring in the project Neo4j, commit a05a8c5, 2: a suggestion produced by static analysis tool, 3: a hallucination suggested by an LLM, 4: a not useful suggestion produced by LLM.
@@
We illustrate the challenges of suggesting code fragments to extract into new methods in a way that humans find useful, as well as the process of unleashing the full potential of LLMs for Extract Method refactoring using a real-world example. Figure 1 shows an Extract Method refactoring performed by Neo4J developers. They extracted statements in lines 157 and 158 into a new method named emptyPropertyArray ( 1 in Figure 1). This refactoring decision is intuitively sound as these two statements collectively fulfill a single responsibility, namely, the creation of an empty array.
@@
When replicating this real-world scenario, existing techniques (Maruyama, 2001; Tsantalis and Chatzigeorgiou, 2011; Charalampidou et al., 2017) for suggesting Extract Method would recommend extracting lines 157–163 ( 2 in Figure 1). This recommendation arises from the fact that the variable values is used later in the code following the two aforementioned statements. However, this selection of statements does not align with the developers’ actually performed refactoring ( 1 in Figure 1) despite the fact that the extraction of lines 157-163 adheres to the static analysis principles (e.g., extracting a program slice) and software-quality metrics (e.g., improving the cohesion of statements in a method) governing existing tools. This finding is in line with research that shows that code-quality metrics do not necessarily capture code quality improvements as perceived by the developers (Fakhoury et al., 2019; Scalabrino et al., 2019).
Pre-trained Large Language Models (LLMs) signify a pivotal shift in the landscape of natural language processing. These models undergo rigorous training on extensive and diverse datasets, including terabytes of texts. In addition to books, articles, and texts from the web, modern datasets also comprise source code at GitHub scale (Kocetkov et al., 2022; Manh et al., 2023). Training on a broad range of texts leads to models that are capable in a wide range of language and code processing tasks (OpenAI, 2023; Brown et al., 2020), all orchestrated through prompt engineering (Reynolds and McDonell, 2021; Wei et al., 2022). Prompt engineering involves the skillful formulation of natural language prompts to direct LLMs’ behavior and responses to solve the task at hand.
Moreover, LLMs stand out as formidable instruments, particularly in the domains of source code generation and refinement. Their ability is evident in their capacity to suggest code improvements, including the detection and rectification of “code smells” (Fowler, 1999; Azeem et al., 2019). Since LLMs are trained on vast collections of real-world code, they can capture common patterns that humans use when writing code. We hypothesize that LLMs can effectively suggest which statements should form new methods based on the patterns of which statements commonly go together. This can narrow the gap between automated suggestions and human intuition.
@@
We revisit the refactoring scenario, only this time we involve an LLM, specifically opting for GPT-3.5 (Brown et al., 2020) as our choice, developed by OpenAI. We provided the LLM with a prompt that included the code snippet in the host method entityGetProperties, followed by a description of the Extract Method refactoring process and an instruction to generate refactoring suggestions for the method. As LLMs produce results that are non-deterministic, we repeated the same prompting and collected suggestions until we reached a fix-point, i.e., the LLM no longer produced new suggestions. In this case, we reached the fix-point in 9 iterations.
@@
The LLM came up with 9 distinct suggestions for which statements to extract, among these including the one that the Neo4J developer performed. In addition to 3 applicable suggestions, the LLM also proposed to extract lines 162–164 ( 3 in Figure 1). If we did so, the code would not even compile because the suggestion starts inside the while loop and goes past the closing bracket. We call these suggestions invalid. For our scenario, the LLM produced 3 invalid suggestions. While the LLM might have seen those code statements being collocated and associated in real-world codebases, because the LLM does not fully understand the semantics of Extract Method and syntactical rules of well-formed code, those associations would render the extracted method not compilable.
@@
Moreover, the LLM suggested extracting the code fragment in line 153 ( 4 in Figure 1), which is a one-liner. While it is possible to extract this line correctly, we consider it not useful as it is improbable for developers to undertake this suggestion. Similarly, the LLM suggested to extract the whole method body into a new method, which does not provide any value for the developers. We call such suggestions not useful. In this experiment, the LLM produced 3 not useful suggestions. The invalid suggestions, together with the not useful ones, represent hallucinations of the LLM.
@@
This experiment showcases the strengths of LLMs: being prolific in generating 9 suggestions, among those is 1 suggestion that aligns with the way how the developers actually performed the refactoring. However, it also showcases the LLM’s weakness: it produced 6 hallucinations, of which 3 are invalid suggestions and 3 are not useful. This shows that it is imperative to exercise caution and not exclusively rely on LLM-generated suggestions when conducting refactorings assisted by LLMs. Moreover, it shows that while LLMs have a huge potential to come up with refactorings that match human’s acceptance criteria, developers that embark on such a journey need to work hard to tap into such potential: they need to tame LLMs non-determinism by repeated prompting (9 in our example), collect those 9 suggestions, and sift through and discard 6 hallucinations. Moreover, as we show in Section 4.3, developers might need to repeat the whole suite of experiments for different LLM “temperature” settings: these control the level of variability for the solutions produced by the LLM. Furthermore, developers might want to experiment with various LLMs (in Section 4.2 we tried three different LLMs), altogether resulting in a combinatorial explosion of experiments.
@@
This is where our tool, EM-Assist, comes alongside to liberate the developers so they can focus on the creative part: using their expertise to decide which of a small number of high-quality suggestions to accept. Next, we present the workflow of using EM-Assist and its core architecture.
3. Technique

### v2

@@ Abstract
Long methods that encapsulate multiple responsibilities within a single method are challenging to maintain. Choosing which statements to extract into new methods has been the target of many research tools. Despite steady improvements, these tools often fail to generate refactorings that align with developers’ preferences and acceptance criteria. Given that Large Language Models (LLMs) have been trained on large code corpora, if we harness their familiarity with the way developers form functions, we could suggest refactorings that developers are likely to accept.
@@
In this paper, we advance the science and practice of refactoring by synergistically combining the insights of LLMs with the power of IDEs to perform Extract Method (EM). Our formative study on 1752 EM scenarios revealed that LLMs are very effective for giving expert suggestions, yet they are unreliable: up to 76.3% of the suggestions are hallucinations. We designed a novel approach that removes hallucinations from the candidates suggested by LLMs, then further enhances and ranks suggestions based on static analysis techniques from program slicing, and finally leverages the IDE to execute refactorings correctly. We implemented this approach in an IntelliJ IDEA plugin called EM-Assist. We empirically evaluated EM-Assist on a diverse corpus that replicates 1752 actual refactorings from open-source projects. We found that EM-Assist outperforms previous state of the art tools: EM-Assist suggests the developer-performed refactoring in 53.4% of cases, improving over the recall rate of 39.4% for previous best-in-class tools. Furthermore, we conducted firehouse surveys with 16 industrial developers and suggested refactorings on their recent commits. 81.3% of them agreed with the recommendations provided by EM-Assist. This shows the usefulness of our approach and ushers us into a new era when LLMs become effective AI assistants for refactoring.
I Introduction
@@
Excessively long methods that encapsulate multiple responsibilities within a single method are challenging to comprehend, debug, reuse, and maintain [1, 2, 3]. To mitigate these issues, software developers use Extract Method (EM) refactoring – a hallmark refactoring [3] supported by all modern IDEs. This refactoring involves moving a block of statements from a host method to a brand new method, passing the used variables as parameters to the new method, and adding a call to the new method from the host method. This refactoring is reported [4, 5, 6] as being among the top-5 most frequently performed in practice, both for manual and automated refactoring.
@@
The process of performing an EM consists of two phases: (i) choosing the statements to extract and (ii) applying the mechanics to perform this refactoring. While the application part has been a staple feature of all modern IDEs, they leave it up to developers to choose the statements to extract. Researchers developed several tools that suggest which statements to extract based on static analysis [7, 1, 8, 9, 10], or machine learning (ML) models [11, 12]. These tools use software quality metrics (e.g., coupling, cohesion), which they optimize based on heuristics or by training ML classifiers.
@@
While these tools adhere to software engineering principles like the Single Responsibility Principle [13], they fail to align with real-world EM instances. We posit that the decision to refactor is both a science and an art. Developers use both their knowledge of software engineering principles and rely on their own experience and subjective interpretation of the code context and what makes a good method. This might explain why developers are reluctant to use automated refactorings [14] and the large gap between high metrics scores and their low acceptance by developers [15, 16, 17, 18, 19].
@@
Recently, Large Language Models (LLMs) [20, 21, 22] are emerging as powerful companions for several software engineering tasks [23, 20, 24, 25, 26]. Given LLMs’ training on extensive code repositories that contain billions of methods written by actual developers, we hypothesize that they are more likely to imitate human behavior by mimicking how developers form functions, thus are likely to suggest refactorings that developers would accept. Our formative study on 1752 EM scenarios from 12 open-source projects revealed that LLMs are very effective in giving expert suggestions. Moreover, they are very prolific, producing on average 27 suggestions per method. However, we also discovered that LLMs are unreliable: 76.3% of their suggestions are hallucinations, i.e., they seem plausible at first, but are actually deeply flawed. We found that 57.4% of the suggested refactorings contain code fragments that are invalid to extract (e.g., would produce compile errors), and 18.9% are not useful (e.g., suggest to extract the whole host method).
To advance the state of the art and practice for refactoring, we bridge several important gaps. First, we bridge the gap between the suggestion of refactorings and developers’ actual practices by grouping statements into methods that resemble human-written code. Second, we bridge the gap between suggesting and applying refactorings by supporting the whole end-to-end process in a way that provides maximum automation while taking into account human input.
@@
We have designed, implemented, and evaluated EM-Assist, an IntelliJ IDEA plugin, that supports Java and Kotlin EM refactorings. EM-Assist synergistically combines (i) the creative capabilities of LLMs, (ii) static analysis techniques to filter and enhance LLM-provided suggestions, and (iii) the full power of a state-of-the-practice commercial IDE to apply refactorings safely. EM-Assist first repeatedly prompts the LLM in a few-shot learning style to generate a diverse range of refactoring suggestions. Subsequently, it eliminates two kinds of hallucinations: (i) it employs static analysis techniques from the IDE to eliminate invalid suggestions (i.e., illegal groupings of code statements), and (ii) it eliminates suggestions that are not useful (e.g., that include the whole body of the host method, or one single line). Then it further enhances the remaining valid suggestions based on program slicing techniques. Following this step, it ranks and prioritizes high-quality suggestions so that it would not overwhelm the developer with too many suggestions. It then presents the ranked suggestions to the developers. Lastly, it encapsulates the user-chosen candidate into a refactoring command and uses the IDE to correctly execute the refactoring.
@@
We designed a comprehensive empirical evaluation to determine the benefits of our novel approach, the pros and cons of using LLMs, and a sensitivity analysis. To determine the effectiveness of EM-Assist, we use two complementary methods. First, we use a publicly available corpus [11] of 122 EMs that other researchers used in the past. The results show that our EM-Assist outperforms state-of-the-art static-analysis tools such as JDeodorant [1], JExtract [27], SEMI [8], LiveRef  [28, 29], and also outperforms ML-based techniques like REMS [12] and GEMS [11]. EM-Assist correctly suggests the ground truth EM refactoring among the top-5 candidates 60.6% of times, compared to 54.2% reported by existing ML models, and 52.2% reported by existing static analysis tools. Second, we significantly expanded the previous community corpus of 122 EMs with a diverse corpus of 1752 actual EM instances from open-source projects. EM-Assist correctly suggests the actual refactoring performed by the developers in 53.4% of cases, an improvement over the 39.4% rate of the previous best-in-class static analysis tool.
@@
Moreover, to assess whether refactoring suggestions generated by EM-Assist are useful, we employed firehouse surveys [30] with 16 industrial developers from a reputable software company. We presented them with refactoring suggestions for lengthy methods they previously committed into their software repository. 81.3% of respondents agreed that the suggestions were beneficial and suitable for application in their codebase. They wish to see EM-Assist in production in IntelliJ IDEA and use it in their daily coding. This demonstrates that EM-Assist generates suggestions that are highly likely to be embraced by developers and ushers us into a new era of refactoring when developers are accepting AI assistants.
In summary, this paper makes the following contributions:
@@
    We present the first approach that automates the full refactoring lifecycle by synergistically combining LLMs and IDEs. Thus, we effectively bridge the gap between refactoring recommendation and application, and we shrink the gap between refactoring suggestions and developer practices.
@@
    We discovered best practices for prompting and tuning LLMs to produce effective refactoring suggestions and we reveal LLMs’ strengths and weaknesses for refactoring.
@@
    Our comprehensive empirical evaluation on a corpus used by others, as well as significantly extended corpus that replicates 1752 real-world refactorings shows that EM-Assist outperforms static analysis-based techniques and ML approaches that suggest EM refactoring. Moreover, our survey with 16 industrial developers provides insights about the reasons why developers accept or reject EM-Assist’s suggestions.
@@
To aid reproducibility, we make EM-Assist, a demo, and all our experimental data available anonymously [31].
II Motivating Example
@@
Figure 1: The numbered code snippets represent (1) an extract function refactoring in the project Neo4j, commit a05a8c5, (2) a suggestion made by static analysis tool, (3) an invalid suggestion from LLM, (4) a not useful suggestion from LLM.
@@
We illustrate the challenges of suggesting code fragments to extract into new methods in a way that aligns with developer preferences, as well as the process of unleashing the full potential of LLMs for EM refactoring. Figure 1 shows an EM refactoring performed by Neo4J developers. They extracted statements in lines 157 and 158 into a new method named emptyPropertyArray ( 1 in Figure 1). This refactoring decision is intuitively sound as these two statements collectively fulfill a single responsibility, namely, the creation of an empty array.
@@
When replicating this real-world scenario, existing techniques [7, 1, 8] for suggesting EM would recommend extracting lines 157–163 ( 2 in Figure 1). This recommendation arises from the fact that the variable values is used later in the code following the two aforementioned statements. However, this selection of statements does not align with the developers’ actually performed refactoring ( 1 in Figure 1) despite the fact that the extraction of lines 157-163 adheres to the static analysis principles (e.g., extracting a program slice) and software-quality metrics (e.g., improving the cohesion of statements in a method) governing existing tools. This finding is in line with research that shows that code-quality metrics do not necessarily capture code quality improvements as perceived by the developers [15, 16].
@@
Next, we used a LLM to suggest EM refactorings for this method, specifically opting for GPT-3.5 [32] because it is cost-effective. We provided the LLM with a prompt that included the code snippet in the host method entityGetProperties, followed by a description of the EM refactoring process and an instruction to generate refactoring suggestions for the method. As LLMs produce results that are non-deterministic, we repeated the same prompting and collected suggestions until we reached a fix-point, i.e., the LLM no longer produced new suggestions. We reached the fix-point in 9 iterations.
@@
The LLM came up with 9 distinct suggestions for which statements to extract, among these including the one that the Neo4J developer performed. In addition to 3 applicable suggestions, the LLM also proposed to extract lines 162–164 ( 3 in Figure 1). If we did so, the code would not even compile because the suggestion starts inside the while loop and goes past the closing bracket. We call these suggestions invalid. The LLM produced 3 invalid suggestions. While the LLM might have seen those code statements being associated together in codebases, because the LLM does not fully understand the semantics of EM and syntactical rules of well-formed code, those associations render the extracted method not compilable.
@@
Moreover, the LLM suggested extracting the code fragment in line 153 ( 4 in Figure 1), which is a one-liner. While it is possible to extract this line correctly, we consider it not-useful as it is improbable for developers to undertake this suggestion. Similarly, the LLM suggested to extract the whole method body into a new method, which does not provide any value for the developers. We call such suggestions not-useful. In this experiment, the LLM produced 3 not-useful suggestions. The invalid suggestions, together with the not-useful ones, represent hallucinations of the LLM.
@@
This experiment showcases the strengths of LLMs: being prolific in generating 9 suggestions, among those is 1 suggestion that aligns with the way how the developers actually performed the refactoring. However, it also showcases the LLM’s weakness: it produced 6 hallucinations, of which 3 are invalid suggestions and 3 are not useful. This shows that it is imperative to exercise caution and not exclusively rely on LLM-generated suggestions when conducting refactorings assisted by LLMs. Moreover, it shows that while LLMs have a huge potential to come up with refactorings that match human’s acceptance criteria, developers that embark on such a journey need to work hard to tap into such potential: they need to tame LLMs non-determinism by repeated prompting (9 in our example), collect those 9 suggestions, and sift through and discard 6 hallucinations. Moreover, as we show in Section IV-D, developers might need to repeat the whole suite of experiments for different LLM “temperature” settings: these control the level of variability for the solutions produced by the LLM. Furthermore, developers might want to use various LLMs (in Section IV-C we tried three LLMs), altogether resulting in a combinatorial explosion of experiments.
@@
This is where our tool, EM-Assist, liberates the developers so they can focus on the creative part: using their expertise to examine a small number of high-quality suggestions.
III Technique